ASseLs g81 63.37 le ou 2 STC 6 ong tel dolel 2 2그 g2SOS 2 228 su그 su iit and Li

All numbers are in million.

Depreciation expense is 1399. Could you please calculate CFO and then FCFE?

The formula for FCFE is CFO - CAPEX * Debt/Assets Ratio.
Answer #1

1. CFO= NI+ NONCASH EXPENSES+ changes in working capital

Find out net income by subtracting ending reserves and initial and do the same for the rest of the components.

NI= 3158.43

+ DEP= 1399

- A/C RECEIVABLE= 114.78

- INVENTORY= 2136.35

+ a/c payable= 1023.61

+ notes payable = 84.11

Thus CF0= 3414.02

Now FCFE= CFO- CAPEX + NET BORROWING

= 3414.02- 2039.7=1374.32

2039 IS THE INCREASE IN NET PPE FROM 2019 T 2020 SO IT IS A CAPEX